OpenStudy (asad786):

WILL MEDAL What is the value of the expression –12 – 3 + 9? A. –6 B. 22 C. 0 D. –15

OpenStudy (asad786):

I THINK A

pooja195 (pooja195):

That's right

OpenStudy (asad786):

I MEAN C

pooja195 (pooja195):

−12−3+9 =−15+9 =−6

OpenStudy (asad786):

add comes before subtract so 9+3=12

pooja195 (pooja195):

You were right with A .-. Why do you think it's C? Don't use PEMDAS here use the left to right rule

OpenStudy (asad786):

oh ok